OpenStudy (anonymous):

rewrite the formula to solve for the amount of money thay will make. 5.5=m/n

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

...

OpenStudy (anonymous):

what is m and n?

OpenStudy (anonymous):

\[5.5 = \frac{ m }{ n }\] now multiply both sides by n

OpenStudy (anonymous):

thats how it already is thay want it a diff way

OpenStudy (anonymous):

what way?

OpenStudy (anonymous):

read my last line

OpenStudy (anonymous):

the way u just put it

OpenStudy (anonymous):

yea do what i told you multiply both sides of the equations by n \[5.5 \times n = \frac{ m }{ n } \times n\]

OpenStudy (anonymous):

now simplify this

OpenStudy (anonymous):

thx i got it

OpenStudy (anonymous):

yw